### v2

@@ 4.4.2 Quadratic Costs
With this cost function we have $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\lambda\cdot\mu}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\lambda}=\frac{c}{2}\mu^{2}\left(\lambda-1\right)^{2}$ for all $\lambda>1$ as desired. In the left panel of Figure 1, the cost function is plotted for $\mu=1$ and $c=2$. One property of it is that $\lim_{\tilde{\mu}\rightarrow\infty}C(\tilde{\mu},\mu)=\lim_{\tilde{\mu}\rightarrow 0}C(\tilde{\mu},\mu)=\infty$, and therefore we have $\tilde{\mu}\in(0,\infty)$ for any $s$.
@@
$W(s,\mathbf{p},\tilde{\mu})=AU(s,\mathbf{p},\tilde{\mu})-C(\tilde{\mu},\mu).$
As before, focus on $p_{1}\neq p_{2}$. When the signal is uninformative, $s=0$, any distortion only causes cost, but does not change $AU(s,\mathbf{p},\tilde{\mu})$. Hence, in this case, $\tilde{\mu}^{*}=\mu$. But what happens when signals are very informative? Suppose $s\rightarrow\infty$. Then any distortion except $\tilde{\mu}=0$ will move the belief about $\omega$ to $\pi=1$. But the cost of choosing $\tilde{\mu}=0$ is infinite, and hence this cannot be optimal. Therefore, it is optimal to not distort at all. The same is true for $s\rightarrow-\infty$. For any other signal realization, $s\in\left(-\infty,0\right)\cup\left(0,\infty\right)$, the direction of the optimal distortion is the same as in Lemma 1. Moreover, $W(s,\mathbf{p},\tilde{\mu})$ is a continuous and smooth function of both $s$ and $\tilde{\mu}$, and hence also $\tilde{\mu}$ is a continuous and smooth function of $s$. To get a better intuition of how $\tilde{\mu}^{*}$ changes with $s$, I plotted it for $\mu=1$ and different values of $c$ in the right panel of Figure 1. The next lemma states the above intuitions formally:
Lemma 4.
Suppose $\Delta>1$. If $s\in\{-\infty,0,\infty\}$, then $\tilde{\mu}(s)=\mu$. Otherwise, and if $s>0$, then $\tilde{\mu}(s)<\mu$, and there exists $\bar{s}>0$ such that $\tilde{\mu}(s)$ decreases in $s$ if $s\in[0,s^{+})$, and it increases in $s$ if $s>s^{+}$. If $s<0$, then $\tilde{\mu}(s)>\mu$, and there exists $s^{-}<0$ such that $\tilde{\mu}(s)$ decreases in $s$ if $s\in(s^{-},0]$, and it increases in $s$ if $s<s^{-}$. Moreover, for every $s$, the absolute distortion $\left|\mu-\tilde{\mu}\right|$ decreases in $c$
@@
Now let us discuss the implications of the lemma for voters’ decisions at the ballot, and by extension, for candidates’ incentives to offer the optimal policy. First, just as in the case of fixed costs, if $s<0$, voters tend to amplify the signals, leading more voters to vote informatively compared to the absence of motivated beliefs. This improves the ability to select the right policy through an election. Therefore, to understand the conditions under which candidates ignore information and choose policy 0 regardless of $\omega$, we need to focus on voters receiving $s>0$. These voters interpret signals as less informative than they truly are, resulting in fewer voters casting informative votes compared to the absence of motivated beliefs. Because $q<\tilde{\pi}$, if either the distortion of $\mu$ is significant or the signal is too weak, such a voter may cast a ballot for policy 0 despite $s>0$. Intuitively, as $c$ increases, voters choose to distort $\mu$ less and less. Consequently, $\lim_{c\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$. In this case, by Assumption 2, a majority will always cast a ballot for $p=\omega$, giving candidates an incentive to heed their information. However, if $c$ is small, we revert to the result from Proposition 3. By continuity, there exists a $\hat{c}>0$ such that if $c\in[0,\hat{c}]$, motivated beliefs will lead to distortions and voting behavior that prevent candidates from following their information.
Proposition 6.
Suppose $\Delta>1$. There exists $\hat{c}>0$ such that iff $c<\hat{c}$, then there exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, where $\tilde{\mu}^{*}$ follows from Lemma 4, in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
4.4.3 Comparing Fixed and Continuous Costs
The analysis in this section has given us more nuanced results for the optimal distortion of beliefs. Moreover, in both extensions discussed so far, the inactive equilibrium prevails as the unique equilibrium if the cost of distorting informativeness of signals is low enough. However, the two models discussed differ significantly in how voters choose to distort information and hence beliefs, see Figure 2, allowing to derive testable predictions.
Figure 2: Equilibrium belief distortion (solid, red curve) with fixed costs (left panel) and with a strictly convex cost function (right panel). The blue and dashed curve is the Bayesian belief.
With fixed costs, only the worst signals are distorted, both in the case of good news ($s<0$) and bad news ($s>0$). When news are bad, the worst signals are the ones suggesting $\omega=1$ the strongest. This leads to low anticipatory utility and thus the signal is interpreted as not informative. However, if bad news is weak, then the belief does not change a lot anyway, and hence choosing $\tilde{\mu}\neq\mu$ at a cost is not attractive. When news is good, the worst signals are the ones suggesting $\omega=0$ the least. Weak good news leads to little updating of beliefs, and overreacting to information is better for anticipatory utility, even if it is somewhat costly. However, if good news is strong enough, then overreacting to the signal does not change beliefs that much anymore, and hence distorting beliefs is not in a voters interest.
If, instead, we consider a continuous and convex cost function, the results are to some degree reversed. In fact, when $s$ is very large, implying a voter receives strong bad news, then interpreting the signal as not informative is very costly, and hence a voters will be better served simply interpreting the signal in an (almost) Bayesian fashion. The same is true when $s$ is close to zero. However, bad news of intermediate strength will be downplayed significantly. When a voter receives good news we find similar results. If $s$ is very low, the voter’s belief about $\omega$ is close to zero even if $\tilde{\mu}=\mu$, and hence distorting $\mu$ is not very attractive. The same is true if good news is very weak. However, if good news is of intermediate strength, then distorting $\mu$ upwards is attractive.
5 Discussion and Some Evidence
The formal analysis highlights at least two important determinants of the efficiency of policy-making in the face of climate change, rhetoric and trust in government.
The Importance of Rhetoric:
When $\Delta>1$, there is no equilibrium in which office-motivated candidates campaign with the policies needed for severe climate change, even if they know for sure that $\omega=1$. Of course, in real life, we do not discuss the issue of climate change in terms of baseline welfare losses, but rather using verbal descriptions. In recent years, an increasing number of people have stopped using the relatively neutral terms “global warming” or “climate change” and have instead adopted more extreme and loaded language such as “climate crisis” or “climate emergency.” Indeed, a comparison of Google search trends in the U.S. from January 2014 to July 2024 shows that, starting around 2019, the relative frequency with which these terms are used has been increasing, see Figure 3. One reason for this change is that many activist groups started to use a more aggressive rhetoric in an effort to emphasize the importance of swift climate policy changes. For example, the activist group Extinction Rebellion writes on their homepage: [^3] See https://rebellion.global/ (accessed on 24/09/2024).
(a) Climate Change
(b) Climate Emergency
(c) Climate Crisis
Figure 3: Five months moving averages of Google search trends in the U.S. for “Climate Change”, “Climate Emergency”, and “Climate Crisis” from January 2014 to July 2024.
This is an Emergency. Life on Earth is in crisis. Our climate is changing faster than scientists predicted and the stakes are high. Biodiversity loss. Crop failure. Social and ecological collapse. Mass extinction. We are running out of time, and our governments have failed to act.
The model suggests that such a shift in language may have undesired consequences, because it may lead voters to ignore information suggesting that climate change is a problem, whereas information stating the opposite is trusted more than it should. And in fact, the fraction of U.S. Americans who stated in PEW polls that they “personally care a great deal about the issue of climate change” increased between 2016 and 2018 from 36% to 44%, but then dropped to 37% in 2023. At the same time, the fraction of people who stated to either not care too much or to not care at all about the issue first decreased from 26% in 2016 to 22% in 2018, but then increased again to 27% in 2023.444See https://www.pewresearch.org/science/2023/10/25/how-americans-view-future-harms-from-climate-change-in-their-community-and-around-the-u-s/. These changes can be explained by the model, and they create stronger incentives for politicians to ignore their information about the true state of the world when designing policies.
That framing and language may have an important effect on wether people learn or not is well established in psychology. For example, Smith and Mackie (2007) mention the importance of not creating too much anxiety in receivers when the intent is to change their attitudes:
To be successful, fear appeals must arouse just the right amount of anxiety by showing the negative consequences […] that will follow if behavior doesn’t change. The provision of an explicit avenue of action […] eliminates the anxiety by showing how the negative consequences can be avoided.
Smith and Mackie (2007), page 257
Excessively negative rhetoric may create too much anxiety and may hence prevent attitude changes and learning, see also McDonald, Chai, and Newell (2015). The risk to trigger such excessive emotional responses is particularly large when people’s general livelihood seems threatened, which is true in the case of climate change. Writing explicitly about the issue climate change, Clayton, Manning, Krygsman, and Speiser (2017) state that
[t]he ability to process information and make decisions without being disabled by extreme emotional responses is threatened by climate change. Some emotional response is normal, and even negative emotions are a necessary part of a fulfilling life. In the extreme case, however, they can interfere with our ability to think rationally, plan our behavior, and consider alternative actions.
Clayton, Manning, Krygsman, and Speiser (2017), page 16
Other papers that highlight the importance of framing for the efficacy to learn about climate change include Spence and Pidgeon (2010), Jones and Peterson (2017), and Ngo, Poortvliet, and Klerkx (2022). Hence, language or rhetoric may indeed be important factors in explaining voters’ attitudes towards climate change, and therefore may as well be determinants of political inactivity in the face of climate change.
The Importance of Trust in Government:
On the other hand, Trust in the Government to choose appropriate policies for the problems at hand matters. In the model this is captured by $\kappa(s,\mathbf{p},\tilde{\mu})$. If voters belief the correct policy will be chosen, then $\kappa(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\tilde{\mu})$, and in this case an efficient equilibrium exists if $\Delta<1$. However, if voters are pessimistic about politician’s ability or incentives to implement the right policies—$\kappa(s,\mathbf{p},\tilde{\mu})$ is then independent of the signal $s$—then this becomes self-confirming, leading politicians to ignore their information, and the optimal policy may not be chosen. Hence, the model suggests that there is a positive relationship between efforts against climate change and trust in government. Note that in a model with no motivated beliefs, such a prediction could not emerge, because the belief about climate change is independent of trust in government, and hence so is the implemented policy. It is the interaction of trust in government $\kappa$ and motivated beliefs $\tilde{\mu}$ that causes this result.
Of course, trust in government is not the only determinant of progress against climate change. Different countries face different costs of adopting new policies. In the model, this is captured by $\beta$. When $\beta$ is large, inaction may be very costly, while if $\beta$ is small, inaction has little consequences. Empirically, $\beta$ could be interpreted as economic cost of (not) dealing with climate change in an appropriate way. Often, anti-climate change policies aim at reducing CO2 emissions. This may be quite costly for countries rich in oil and gas, while it is less costly for other countries. Hence, countries rich in fossil energy sources are less likely to make progress against climate change than countries with little or no such resources.
Statistic N Mean St. Dev. Min Max
CCPI 19 53.66 13.59 26.47 79.61
TGI 19 51.33 13.07 25.63 77.54
Oil 19 9.95 38.82 0.00 170.00
Risk 19 2.20 1.15 1.00 5.40
Table 1: Summary statistics of included variables.
We can test the hypotheses using cross-sectional data. As the dependant variable, I use the Climate Change Performance Index (CCPI) published jointly by Germanwatch, the NewClimate Institute, and the Climate Action Network International. [^5] See https://ccpi.org/. The CCPI takes values between 0 and 100, where higher values indicate that a country performs better in enacting appropriate policies against climate change. The considered policies include reductions in greenhouse gas emissions (40%), expansion of renewable energy sources (20%), general energy use (20%), and other climate policies (20%). As controls for trust in government I use the OECD’s Trust in Government Indicator (TGI). [^6] See https://www.oecd.org/en/data/indicators/trust-in-government.html The TGI also takes values between 0 and 100 and equals the share of people reporting to have confidence in the national government. As additional control variables, I use countries’ confirmed oil reserves in billion barrels (Oil), which can be interpreted as a measure of the cost of climate change policies, and the Global Climate Risk Index (Risk), published by Germanwatch, which can be interpreted as a measure of the benefits of climate change policies. [^7] See https://www.germanwatch.org/en/cri In Table 1 I summarize the included variables.
Figure 4: Scatterplot of Climate Change Performance Index (CCPI) in 2023 and the OECD Trust in Government Indicator (TGI) from 2022 with OLS trendline.
Unfortunately, the number of countries for which all of these variables are available is small ($N=19$), and hence power is limited. Nevertheless, let us now take a first look at the data. Figure 4 shows a scatter plot of CCPI and TGI for the different countries including a trend line. The plot indicates a positive relation between TGI and CCPI, as predicted by the model. The Scandinavian countries have the highest trust in government and also show some of the highest CCPI scores. The US, Poland, or Colombia have relatively low trust in government, and they also have low CCPI scores. Canada has a low CCPI score despite having a medium-high TGI score. [^8] Note that the theoretical analysis is valid only for democracies. However, because all countries in the sample are classified either as Democracy or Full Democracy in the Polity IV data series, and because the sample size is very small, I do not control for polity scores in the regressions. However, Canada has significant confirmed Oil reserves.
To shed further light on the empirical relation between trust in government and climate change policies, I regressed CCPI on TGI and a subset of the other control variables. In particular, I estimate the following regression equation:
$CCPI_{i}=\alpha+\beta\,TGI_{i}+\gamma\,\mathbf{X}_{i}+\epsilon_{i}$
The estimate of $\beta$ captures the relationship we are interested in. Table 2 below shows the regression results. Despite the low power, $\hat{\beta}$ is positive and statistically significant in all specifications. Moreover, a one standard deviation increase of TGI leads to an increase of CCPI between 0.463 and 0.495 standard deviations, depending on the chosen specification. Therefore, the relation appears to be of relevance. Checking for potential heteroscedasticity, the $p$-values of the Breusch-Pagan test are larger than 0.4 for all specifications, suggesting that heteroscedasticity is no problem. Inspecting the Q-Q plots of the different specifications (see Figure 5 in Appendix B), further supports this conclusion. Hence, there indeed seems to be a significant and positive correlation between CCPI and TGI, as suggested by the model.
Dependent variable:
CCPI
(1) (2) (3) (4)
TGI $0.484^{**}$ $0.481^{**}$ $0.498^{*}$ $0.515^{**}$
$(0.223)$ $(0.193)$ $(0.250)$ $(0.216)$
Oil $-0.168^{**}$ $-0.170^{**}$
$(0.065)$ $(0.067)$
Risk2022 $0.397$ $0.982$
$(2.832)$ $(2.458)$
Constant $28.825^{**}$ $30.629^{***}$ $27.249$ $26.754^{*}$
$(11.803)$ $(10.251)$ $(16.566)$ $(14.316)$
Observations 19 19 19 19
R2 0.217 0.446 0.218 0.452
Adjusted R2 0.170 0.377 0.120 0.343
F Statistic 4.699∗∗ (df = 1; 17) 6.451∗∗∗ (df = 2; 16) 2.224 (df = 2; 16) 4.128∗∗ (df = 3; 15)
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
Table 2: Regression Results.
@@ 6 Conclusion
This paper illustrates the negative impact that voter beliefs and expectations can have on political decision-making, particularly in the context of climate change policy. The model demonstrates that when voters are inclined to dismiss the severity of climate change, this skews the equilibrium toward suboptimal policy choices by incentivizing office-motivated politicians to act in ways that contradict their own information. The analysis reveals that—in the presence of motivated reasoning—trust in the functioning of government plays a critical role for the efficiency of policy-making. A first look at the data confirms this result. Moreover, the paper suggests that political rhetoric plays an important role as well: if politicians, scientists, journalists, and activists aim to contribute to better climate change policy, they must navigate a delicate balance between conveying information truthfully and avoiding the risk of pushing voters into denial.
@@ A.1 Proof of Lemma 1
$\begin{array}[]{rcl}\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}&>&0\\ \Leftrightarrow-\dfrac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}&>&0\\ \Leftrightarrow s\left[\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)\right]&<&0\end{array}$
Here we have $s$ and the expression in brackets that jointly determine the derivative’s sign. Consider next the term in parentheses:
$\beta+\Delta-\kappa(s,\mathbf{p},\tilde{\mu})(1+\beta)>0\Leftrightarrow\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}\equiv\frac{\Delta+\beta}{1+\beta}$
If $\Delta>1$, then $\tilde{\kappa}>1$, and hence we cannot have $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$. It follows that in this case $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$. Hence, $\tilde{\mu}^{*}=\infty$ if $s>0$ and $\tilde{\mu}^{*}=0$ if $s<0$. If $s=0$, any $\tilde{\mu}$ yields the same $AU$ and hence $\tilde{\mu}^{*}=\sigma$.
If $\Delta<1$, $\tilde{\kappa}<1$ as well. Hence, if $\kappa(s,\mathbf{p},\tilde{\mu})<\tilde{\kappa}$, then again $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$. It follows that $\tilde{\mu}^{*}=\infty$ if $s>0$, $\tilde{\mu}^{*}=0$ if $s<0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$. To the contrary, if $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$. Hence, $\tilde{\mu}^{*}=0$ if $s>0$, $\tilde{\mu}^{*}=\infty$ if $s<0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$. If $\kappa(s,\mathbf{p},\tilde{\mu})=\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\sigma})}{\partial\tilde{\mu}}\right]=0$, and any $\tilde{\mu}$ yields the same anticipatory utility. Thus, $\tilde{\mu}^{*}=\sigma$.
Finally, if $\Delta=1$, then we cannot have $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$. Thus, $AU(s,\mathbf{p},\tilde{\mu})$ is either flat in $\tilde{\mu}$ if $\kappa(s,\mathbf{p},\tilde{\mu})=1$, implying any $\tilde{\mu}$ yields the same anticipatory utility, and we have $\tilde{\mu}^{*}=\sigma$, or the situation resembles the case with $\Delta>1$ discussed above.∎

### v5

@@ 5.1.2 Cognitive Costs
With this cost function we have $\left.C^{C}(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\delta\cdot\mu}=\left.C^{C}(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\delta}=\frac{c}{2}\mu^{2}\left(\delta-1\right)^{2}$ for all $\delta>1$. In the left panel of Figure 2, the cost function is plotted for $\mu=1$ and $c=2$. One property of it is that $\lim_{\tilde{\mu}\rightarrow\infty}C^{C}(\tilde{\mu},\mu)=\lim_{\tilde{\mu}\rightarrow 0}C^{C}(\tilde{\mu},\mu)=\infty$, and therefore we have $\tilde{\mu}\in(0,\infty)$ for any $s$.
@@
$W^{CC}=AU(s,\mathbf{p},\tilde{\mu})-C^{C}(\tilde{\mu},\mu).$
[…]
@@
If $s<0$, voters choose optimally at the ballot despite $\tilde{\mu}\neq\mu$. For $s>0$, this is not necessarily the case. These voters interpret signals as less informative than they truly are and may therefore cast a ballot for policy 0 despite $s>0$. Intuitively, as $c$ increases, voters choose to distort $\mu$ less. Consequently, $\lim_{c\to\infty}\tilde{\mu}^{*}(s)=\mu$. In this case, a majority will always cast a ballot for $p=\omega$, giving candidates an incentive to heed their information. However, if $c$ is small, the outcome reverts to the result in Proposition 3. By continuity, there exists a $\hat{c}>0$ such that if $c\in[0,\hat{c}$], motivated beliefs lead to voting behavior that prevents candidates from following their information. Conversely, if $c>\hat{c}$, then despite choosing $\tilde{\mu}^{*}\neq\mu$, a majority of voters will vote for policy 1 in state 1, thereby creating incentives for candidates to follow their information and campaign on policy 1.
5.2 Policy Motivated Politicians
Suppose candidates not only care about being elected but also value welfare, as captured by (1). Knowing which state has materialized, candidates are aware of which policy maximizes welfare. Thus, one might hypothesize that adding a policy motivation would improve the situation.
However, this is only partially correct. On the one hand, if policy motivation is sufficiently important compared to simply being elected to office, there is indeed an equilibrium in which candidates always choose $\mathbf{p}^{*}=(\omega,\omega)$. To see this, assume the spoils of office have value $V$ and that each candidate receives $V+u(p,\omega)$ if elected and $u(p,\omega)$ otherwise.
$\frac{V}{2}-\Delta\geq V-\Delta-b\Leftrightarrow b\geq\frac{V}{2}.$
If, however, $b<\frac{V}{2}$, there remains a unique pure strategy equilibrium with $\mathbf{p}^{*}=(0,0)$. Moreover, increasing the spoils of office $V$ decreases the chance for an efficient equilibrium to exist.
However, even if the spoils of office are sufficiently low, the inefficient equilibrium continues to exist. Assume candidate 1 expects candidate 2 to choose $p_{2}=0$. In this case, the unique best response is to also choose $p_{1}=0$. The reason is that now policy 0 wins with certainty. Candidate 1 cannot influence the policy, but she can increase her chances of winning. Therefore, despite policy motivations, and even in the absence of any office motivation, there always exists an equilibrium with $\mathbf{p}^{*}=(0,0)$. This logic is closely related to the one in Calvert, (1985), who shows that the median voter theorem may continue to hold even in the presence of policy motivations.
@@ 6 Conclusion
This paper illustrates the negative impact that voter beliefs and expectations can have on political decision-making, particularly in the context of climate change policy. The model demonstrates that when voters are inclined to dismiss the severity of climate change, this skews the equilibrium toward suboptimal policy choices by incentivizing office-motivated politicians to act in ways that contradict their own information. The analysis reveals that—in the presence of motivated reasoning—trust in the functioning of government plays a critical role for the efficiency of policy-making. Moreover, the paper suggests that political rhetoric plays an important role as well: if politicians, scientists, journalists, and activists aim to contribute to better climate change policy, they must navigate a delicate balance between conveying information truthfully and avoiding the risk of pushing voters into denial.
@@ A.1 Proof of Lemma 1
$\begin{array}[]{rcl}\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}>0\Leftrightarrow-\dfrac{2(1-q)q(1-\rho_{0})\rho_{0}(1-\rho_{1})\rho_{1}s\left[b+\Delta-\kappa(1+b)\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{\sigma^{2}\left(q\rho_{1}(1-\rho_{1})+(1-q)\rho_{0}(1-\rho_{0})e^{-\frac{2\tilde{\mu}s}{\sigma^{2}}}\right)^{2}}&>&0\\ \Leftrightarrow s\left[b+\Delta-\kappa(1+b)\right]&<&0\end{array}$
The realization of $s$ and the expression in brackets jointly determine the derivative’s sign. Consider next the term in parentheses:
$b+\Delta-\kappa(1+b)>0\Leftrightarrow\kappa<\tilde{\kappa}\equiv\frac{\Delta+b}{1+b}$
If $\Delta>1$, then $\tilde{\kappa}>1$, and hence we must have $\kappa<\tilde{\kappa}$. It follows that $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$. Hence, $\tilde{\mu}^{*}$$\infty$ if $s<0$ and $\tilde{\mu}^{*}=0$ if $s>0$. If $s=0$, any $\tilde{\mu}$ yields the same $AU$ and hence $\tilde{\mu}^{*}=\sigma$.
If $\Delta<1$, then $\tilde{\kappa}<1$ as well. Hence, if $\kappa<\tilde{\kappa}$, then again $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=-\text{Sign}\left[s\right]$. Thus, $\tilde{\mu}^{*}$$\infty$ if $s<0$, $\tilde{\mu}^{*}=0$ if $s>0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$. To the contrary, if $\kappa>\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=\text{Sign}\left[s\right]$. Hence, $\tilde{\mu}^{*}=0$ if $s<0$, $\tilde{\mu}^{*}$$\infty$ if $s>0$, and $\tilde{\mu}^{*}=\sigma$ if $s=0$. If $\kappa=\tilde{\kappa}$, then $\text{Sign}\left[\dfrac{\partial AU(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right]=0$, and any $\tilde{\mu}$ yields the same anticipatory utility. Thus, $\tilde{\mu}^{*}=\mu$.
[…]